stuck on these two parts,please help.

An aircraft engine takes in an amount 9200 J of he
Answer #1

A.

work done = heat in - heat out

W = 9200 - 6600 = 2600 J

B.

efficiency = work out/heat in

n = 2600/9200 = 0.2826

n% = 28.26 % efficiency